# Cycle Consistency as Reward: Learning Image-Text Alignment without Human Preferences

### v1

@@ Appendix E Additional Results
Figure 11: Best-of-$N$ results on DeCapBench for different metrics. Overall, our model increases the level of detail in captions while avoiding severe hallucinations.
Figure 12: Best-of-$N$ results on T2I-CompBench for different metrics. Optimizing with our reward model generally improves results, while VQAScore excels at following positional relationships.
Figure 13: Relative performance gain on T2I-CompBench from Best-of-$1$ to Best-of-$100$ across 6 categories. We mark CycleReward-T2I’s performance with a dashed line in other charts for comparison. While each metric demonstrates category-specific strengths, the human-supervised ImageReward achieves the most balanced overall performance, followed closely by CycleReward-T2I.
@@ E.2 Best-of-N
We provide Best-of-N results for individual categories of T2I-CompBench [33] in Figure 13. Our metric is most effective for complex prompts, whereas the VQAScore excels in spatial relationships.
@@ E.3 Winoground
E.4 DPO
@@ Appendix F Reward Model Trends
We investigate how text and image properties affect different metrics’ alignment preferences for the following factors: caption density, object hallucination, image density, and resolution in Figure 15. For each specific factor, we plot the alignment score for individual image, text pairs based on the relevant image or text characteristic. The title of each plot reports the Pearson correlation coefficient between the alignment score and respective factor. We also display the line of best fit. Note that the scale and range of alignment scores are different and not comparable between metrics. Because of this we instead focus on overall trends and correlations between each factor and alignment.
@@
To examine which reward models generally prefer long or short captions, we first create a dataset of images paired with captions of various lengths. We utilize the test and validation sets of the DCI [83] dataset for this task, where each image is paired with a long, descripitive text. For each image, we use an LLM (Meta-Llama-3.1-8B-Instruct [19]) to create captions of different lengths but asking for summaries with different numbers of words, similarly to Huh et al. [35]. We ask for summaries of lengths 5, 10, 20, …, 100 words, and sample 5 different captions for each length with temperature 0.6 and top p 0.9. This results in 11241 unique image, caption pairs after eliminating duplicates and removing ”here is a summary” text.
In Figure 15(top row), we plot the alignment trend for different metrics versus caption length. The Pearson correlation coefficient is reported at the top of each plot. Note that the alignment score scales are different for each metrics and therefore not directly comparable. Instead, we focus on the overall trends. Because captions can be informative or contain mistakes regardless of their lengths, we expect these plots to be noisy. All methods, except for CycleReward-I2T and ImageReward, have positive Pearson Correlation coefficients - meaning they in general longer captions are preferred. However the correlation between caption length and alignment is much weaker for VQAScore and CLIP compared to CycleReward-Combo and CycleReward-I2T.
@@
For text-to-image, we examine how images of different resolutions affect alignment with the text. To this end, we gather 100 “upsampled” text descriptions created by prompting GPT-4o[61] to add details to short captions from MSCOCO [51]. Text descriptions are encouraged to be visually informative and no longer than 77 tokens. We use SDXL [72] to generate images for each text description at 512$\times$512 resolution. We resize the images to resolutions 256, 128, 64, 32, 16 and compute alignment at each stage in Figure 15(bottom row). For all metrics, alignment is generally not affected when resizing from 512 to 256 and 128 pixels, and then drops off steeply as the resolution goes from 64 to 16. Note that CycleReward and ImageReward preprocess images to be size while CLIP and VQAScore preprocessing resizes image to 336.

### v2

@@ D.3 Diffusion-DPO
Figure 9: DeCapBench Best-of-NN DCScore breakdown. DeCapBench evaluation is performed with DCScore which combines scores for Comprehensiveness and Non-Hallucination in the left and right plots respectively.
@@ Appendix E Additional Results
Figure 11: Best-of-NN results on DeCapBench for different metrics. Overall, our model increases the level of detail in captions while avoiding severe hallucinations.
Figure 12: Best-of-NN results on T2I-CompBench for different metrics. Optimizing with our reward model generally improves results, while VQAScore excels at following positional relationships.
Figure 13: Relative performance gain on T2I-CompBench from Best-of-11 to Best-of-$100$ across 6 categories. We mark CycleReward-T2I’s performance with a dashed line in all charts for comparison. While each metric has category-specific strengths, human-supervised ImageReward achieves the most balanced overall performance, followed closely by CycleReward-T2I.
@@ E.2 Best-of-N
Figure 9 shows DeCapBench Best-of-N results separated into the Non-Hallucination and Comprehensiveness categories used by DCScore [92] during evaluation. All CycleReward models lead to improvement in both categories, but CycleReward-Combo and CycleReward-I2T select the most comprehensive captions, while VQAScore and CycleReward-T2I yield the best non-hallucination scores. Note other metrics such as VQAScore and CLIP have tradeoffs which sacrifice description for accuracy.
@@
Figure 13 shows Best-of-N results for individual categories in T2I-CompBench [33]. Our metric is most effective for complex prompts, whereas the VQAScore excels at spatial relationships.
@@ E.3 Winoground
E.4 More Ablations
Ablation DetailCaps-4870 GenAI-Bench
Best variant (CR-I2T) 58.02 53.49
MSE loss 41.87 40.57
1K images 52.86 44.39
Without filtering 57.28 51.92
E.5 DPO
@@
E.6 Failure Cases
Although we propose cycle consistency as a self-supervised signal for learning image-text alignment, our method has several limitations. A common source of failure is poor reconstructions which mislead preferences determined by cycle consistency seen in Figure 15. Our method also inherits biases from the underlying models used for reconstructions and and similarity measurements. Stable Diffusion 3 has a 77-token limit which limits consideration of longer texts, and LLaVA-1.5-3B can be prone to hallucinations. DreamSim often favors images with similar foregrounds over backgrounds [22], and SBERT is sensitive to text style. Furthermore, we observe worse text-to-image performance, which may partially stem from dataset differences. HPSv2, PickScore, and ImageReward are trained on prompts from real users often describing artwork, whereas CycleReward is trained on LLM-summarized descriptions for natural images. Moreover, cycle consistency primarily considers preservation of information, while other aspects such as aesthetics or style may also affect human preferences. Future work could address these challenges by improving reconstruction quality, prompt diversity, and applying cycle consistency in different scenarios.
Figure 15: Failure cases. (Left): Despite being faithful to the input text, the right image is rejected as the reconstructed text contains hallucinations inconsistent with the original prompt. (Right): The short caption is preferred over the descriptive caption due to an error in text-to-image generation. Under each caption we show the corresponding reconstructed images.
@@ Appendix F Reward Model Trends
We investigate how text and image properties affect different metrics’ alignment preferences for the following factors: caption density, object hallucination, image density, and resolution in Figure 16. For each specific factor, we plot the alignment score for individual image, text pairs based on the relevant image or text characteristic. The title of each plot reports the Pearson correlation coefficient between the alignment score and respective factor. We also display the line of best fit. Note that the scale and range of alignment scores are different and therefore not directly comparable between metrics. Because of this we instead focus on overall trends and correlations between each factor and alignment.
@@
To examine which reward models generally prefer long or short captions, we first create a dataset of images paired with captions of various lengths. We utilize the test and validation sets of the DCI [83] dataset for this task, where each image is paired with a long, descripitive text. For each image, we use an LLM (Meta-Llama-3.1-8B-Instruct [19]) to create captions of different lengths but asking for summaries with different numbers of words, similarly to Huh et al. [35]. We ask for summaries of lengths 5, 10, 20, …, 100 words, and sample 5 different captions for each length with temperature 0.6 and top p 0.9. This results in 11241 unique image, caption pairs after eliminating duplicates and removing “here is a summary“ text.
In Figure 16 (top row), we plot the alignment trend for different metrics versus caption length. The Pearson correlation coefficient $\rho$ is reported at the top of each plot. Because captions can be informative or contain mistakes regardless of their lengths, we expect these plots to be noisy. All methods, except for CycleReward-T2I and ImageReward, have positive Pearson Correlation coefficients - meaning they in general longer captions are preferred. However the correlation between caption length and alignment is much weaker for VQAScore and CLIP compared to CycleReward-Combo and CycleReward-I2T.
@@
For text-to-image, we examine how images of different resolutions affect alignment with the text. To this end, we gather 100 “upsampled” text descriptions created by prompting GPT-4o[60] to add details to short captions from MSCOCO [50]. Text descriptions are encouraged to be visually informative and no longer than 77 tokens. We use SDXL [72] to generate images for each text description at $\times512$ resolution. We resize the images to resolutions 256, 128, 64, 32, 16 and compute alignment at each stage in Figure 16(bottom row). For all metrics, alignment is generally not affected when resizing from 512 to 256 and 128 pixels, and then drops off steeply as the resolution goes from 64 to 16. Note that CycleReward and ImageReward preprocess images to be size while CLIP and VQAScore preprocessing resizes image to 336.